OpenStudy (anonymous):

whats is3.7b+7=8.1b-19.4

OpenStudy (anonymous):

3.7b +7 = 8.1b -19.4 3.7b +26.4 = 8.1b 26.4 = 11.8b 26.4/11.8 = b b = 2.23728

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

anytime :D

OpenStudy (anonymous):

kk heres another one

OpenStudy (anonymous):

what is 0.7z+1.9+0.1=-o.4z